It took 2 mins.15 sec for a volume of chlorine to effuse through a pinhole. How long will it take the same volume of water vapour to effuse through the same hole under the same condition of temperature and pressure?


1
Expert's answer
2022-02-28T11:05:48-0500

The formula for Graham's law of diffusion or effusion of gases as:

rate 1 / rate 2 = √ (mass 2 / mass 1) ,

where: rate 1 and rate 2 - Rates of effusion or diffusion of Gas 1 and 2, respectively, measured moles per unit time.

Let's assume the effused volume of Chlorine and water vapor is 1 L.

M (Cl2) = 71 g/mol

M (H2O) = 18 g/mol

Rate 1 is 1 L / 75 s = 0.013 L/s

Rate 2 H2O = 0.013 / √ (18 / 71) = 0.026 L / s

Time H2O = Volume / rate = 1 L / 0.026 = 39 s